One of the compounds in the "copper cycle" sequence of reactions is copper (II) hydroxide. It is made from the reaction of Cu(NO3)2 and NaOH, as represented in this balanced chemical equation:

Cu(NO3)2 + 2 NaOH --> Cu(OH)2 + 2 NaNO3

What is the theoretical yield of Cu(OH)2 when 43.0 grams of Cu(NO3)2 react with 45.9 grams of NaOH?
Answer #1

Given equation

Cu(NO3)2 + 2NaOH -> Cu(OH)2 + 2NaNO3

molecular mass of

Cu(NO3)2 - 187.5gm

NaOH - 40.5gm

Cu(OH)2 - 97.5 gm

moles of

Cu(NO3)2 - weight/mass = 43/187.5 =0.229 moles

NaOH - 45.9 / 40.5 =1.133 mole

1 mole of Cu(NO3)2 reacts with 2 mole of NaOH .

so 0.229 moles will react with 2X 0.229 i.e 0.458 mole of NaOH,rest left over will remain in the solution.

well 1 mole of  Cu(NO3)2 gives 1 mole of Cu(OH)2 , so 0.229 Cu(NO3)2 mole will give 0.229 mole of Cu(OH)2.

in terms of gm - 0.229 X molecular mass of Cu(OH)2 - 0.229 X 97.5 = 22.32 gm .

Thanks...

I Hope it Helps.